Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 215     Entry time: 31 May 2023, 23:15
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 10.65 keV, 128 pulses at 1.13 MHz and 16 pusles at 141 kHz.
  • Slope reduced at the start of the shift. Approximately 20% slope after that with the patterns used for data collection today.

Optical laser delivery

  • NA

Achievements/Observations

  • Spectrometer setup and aligned. Collected first data with GOTTHARD II.
  • Measured multiple static samples with Bragg magnifier.

Issues

  • No issues.
ELOG V3.1.4-7c3fd00